OpenStudy (anonymous):

i have a physics question that has to do with density + volume and no one has answered me on the physics subject so can i ask it on here???

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

What is the volume of a 612 g piece of aluminum? The density of aluminum is 2.70 g/cm^3.

OpenStudy (anonymous):

ive been trying to figure it out but i just dont know how :(

OpenStudy (anonymous):

\[D = \frac{ M }{ V }\] density = mass / volume = 2.70 g / cm^3 here you are given the mass, M = 612 g and density, D = 2.70 g / cm^3 use D = M/V with those two values to solve for V

OpenStudy (anonymous):

i dont know :/

OpenStudy (anonymous):

\[2.70 = \frac{ 612 }{ V }\]

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

solve for V in the equation. you will get V = 612 / 2.70

OpenStudy (anonymous):

226 would be V then?